The following cost data relate to the office overhead costs of the Vermillion, Co. during the just-completed year:

Office Rent 25,060
office supplies 5,820
utilities, office 14,800
Salary, office manager 54.750
Salary, receptionist 34,320
Depreciation—office equipment 10,920
Software licensing fees 24,430
Other miscellaneous expenses 30,950
  Total overhead costs 201,050

The company uses a predetermined overhead rate to apply overhead cost to each tax job. The rate for the year was $5 per tax-preparer hour. A total of 41,700 tax preparer hours was recorded for the year.

Compute the amount of under- or overapplied overhead cost for the year.

Overhead   
Answer #1

You need to take care of the sign .. (whether only positive or negative is also required) but this is the case of over-applied overhead.

a Applied overhead = 208500
=41700*5
b Actual overhead = 201,050
c=b-a overapplied overhead = -7,450